What concentration of propanoic acid, HC2H5CO2(aq), is needed to prepare a buffer system that contains 0.066 M sodium propanoate, C2H5CO2Na, and has a pH of 4.29 at 25°C? The Ka value of propanoic acid is 1.3 ✕ 10−5.
Answer #1

pKa = - log ( Ka )

= - log ( 1.3*10^-5)

= 4.886

pH = pKa + log ( C2H5CO2Na / C2H5CO2H)

= 4.29 = 4.886 + log ( 0.066 / C2H5CO2H)

= C2H5CO2H = 0.2603 M

If you have any query please comment

If you satisfied with the solution please like it thankxx